### 25 Apr 2026 v3

# AgentHER: Hindsight Experience Replay for LLM Agent Trajectory Relabeling

Abstract
LLM agents fail on the majority of real-world tasks—GPT-4o succeeds on fewer than 15% of WebArena navigation tasks and below 55% pass@1 on ToolBench (Zhou et al., 2024; Qin et al., 2024)—yet every failed trajectory is routinely discarded, wasting the dominant source of collected experience. We introduce AgentHER, which recovers this lost training signal by adapting the Hindsight Experience Replay principle (HER; Andrychowicz et al., 2017) to natural-language agent trajectories. The key insight is simple: a trajectory that fails goal $A$ is often a correct demonstration for some achievable alternative goal $B$. AgentHER realises this idea through a four-stage pipeline—failure classification, outcome extraction, LLM-guided prompt relabeling with confidence gating, and data packaging—that converts discarded failures into high-quality SFT, DPO, and ShareGPT training data, with both zero-cost rule-based and LLM-judge implementations. On WebArena (Zhou et al., 2024) and ToolBench (Qin et al., 2024), AgentHER improves over success-only SFT by +7.1–11.7 pp across four model families (GPT-4o, Qwen2.5-72B/7B, LLaMA-3.1-8B), while achieving $\mathbf{2\times}$ data efficiency—matching baseline performance with only 50% of successful demonstrations. Gains are consistent from 1.5B to 72B parameters (+5.8–9.2 pp) and compound under iterative redeployment (+2.1 pp over additional rounds). Human evaluation confirms 97.7% relabeling precision under multi-judge verification. Code: https://github.com/alphadl/AgentHER
1 Introduction
Autonomous LLM agents are deployed across web navigation (Zhou et al., 2024), API orchestration (Qin et al., 2024), and simulation (Park et al., 2023; Shen et al., 2023). Despite remarkable capabilities, these agents fail on the majority of tasks: GPT-4o achieves only 14.3% success on WebArena (Zhou et al., 2024), and ToolBench pass@1 remains below 55% (Qin et al., 2024).
[…]
The data-waste problem.
Standard agent-training pipelines discard failures (Figure 1). Given 60–75% failure rates, this wastes the majority of collected data. Failures are not random noise: they are rich, coherent executions that frequently produce correct intermediate results under the wrong goal. An agent tasked with “find copper wire under \$5/kg” that searches seven suppliers and identifies MicroMetals at \$5.30/kg (just above the cap) has performed a complete, correct price comparison—ideal training data for a re-framed goal.
Connection to HER.
Hindsight Experience Replay (Andrychowicz et al., 2017) converts sparse-reward RL failures by substituting the intended goal with the one actually achieved. Lifting HER to LLM agents requires: (i) understanding what was achieved across multi-step tool interactions, and (ii) synthesising a natural-language prompt that the trajectory genuinely satisfies. Both are tasks where capable LLMs excel.
AgentHER.
We propose a four-stage pipeline (Figure 2) that automates hindsight relabeling at scale, extending it with two new mechanisms: failure-severity weighting (downweights trajectories with major reasoning flaws; inspired by MQM-style error analysis) and multi-judge verification (two independent judges must agree before a relabeling is accepted), which together reduce label noise from 5.9% to 2.3%.
Contributions.
  • Novel formulation: a systematic pipeline bridging goal-conditioned HER to natural-language agent data synthesis (as opposed to inference-time memory, e.g. ECHO; Hu et al., 2025), with explicit treatment of goal-text faithfulness and factual grounding that distinguishes it from concurrent hindsight relabeling work (HSL; Li et al., 2026) and prior failure-reuse and goal-relabeling work (see Section 2). Unlike HSL, AgentHER targets exclusively offline-collected failures and relabels the top-level goal with multi-judge verification.
  • Two robustness mechanisms: failure-severity weighting (inspired by MQM-style error analysis) and multi-judge verification are engineering improvements that raise pipeline precision from 94.1% to 97.7% and reduce label noise from 5.9% to 2.3%. They are not claimed as independent contributions but as best-practice design choices that make AgentHER production-ready.
  • Consistent gains: +7.1–11.7 pp over SFT-Success across two benchmarks and four model families, scaling from 1.5B to 72B, with $2\times$ data efficiency and +11.0 pp after three iterative rounds. Results are averaged over 3 seeds (std <<0.5 $p$).
2 Related Work
Hindsight Experience Replay and Goal-Conditioned RL.
HER (Andrychowicz et al., 2017) converts sparse-reward RL failures by substituting the achieved state as the new goal (Kaelbling, 1993; Sutton et al., 1999; Mnih et al., 2015). Lifting HER to LLM agents requires understanding what was achieved across unstructured multi-step tool interactions, and synthesising a natural-language prompt that the trajectory genuinely satisfies—two capabilities absent from prior RL formulations. Existing failure-reuse work in NLP selects or reweights (goal, trajectory) pairs (Peng et al., 2023; Gulcehre et al., 2023; Hosseini et al., 2024) but never generates a new achievable goal; that reverse-engineering step is AgentHER’s core novelty. Concurrently, Hu et al. (2025) propose ECHO (Experience Consolidation via Hindsight Optimization), a prompting framework that adapts HER for online LM agents: the model identifies alternative subgoals from failed trajectories and maintains optimized trajectory descriptions in a scratchpad memory to improve sample efficiency at inference time. AgentHER differs in targeting offline training data augmentation: we relabel only the goal (user prompt) while keeping the trajectory unchanged, and output SFT/DPO datasets for fine-tuning rather than inference-time memory; the two approaches are complementary. Also concurrently, Li et al. (2026) propose Hindsight Supervised Learning (HSL), which similarly uses an auxiliary LLM to relabel agent trajectories in hindsight; HSL mines all goals achieved along a trajectory (including intermediate subgoals during successful runs) and maintains a continuously refreshed on-policy buffer, while adding irrelevant-action masking and demonstration reweighting. AgentHER is complementary: it focuses exclusively on failed trajectories from offline collections, relabels the top-level goal while preserving the full trajectory, introduces multi-judge verification and failure-severity weighting to control label noise, and targets WebArena/ToolBench web-navigation and API-use settings rather than the embodied-agent benchmarks (ALFWorld, PlanCraft, WebShop) studied in HSL.
LLM Agents, Training, and Self-Improvement.
ReAct (Yao et al., 2023b; a) and Toolformer (Schick et al., 2023) established reasoning-and-action agents evaluated on benchmarks with high failure rates (Zhou et al., 2024; Qin et al., 2024; Liu et al., 2024). FireAct (Chen et al., 2023) and AgentTuning (Zeng et al., 2024) fine-tune on curated successes; Reflexion (Shinn et al., 2023) and ExpeL (Zhao et al., 2024) improve via online verbal RL and rule extraction; Trial-and-Error (Song et al., 2024) explores mistake-driven recovery. All require successes or live rollouts; AgentHER converts existing failures offline with no extra environment interaction. Self-Instruct (Wang et al., 2023), STaR (Zelikman et al., 2022), and DPO (Rafailov et al., 2023) target single-turn tasks; AgentHER’s multi-judge verification (Gou et al., 2024; Lightman et al., 2024; Du et al., 2024; Liang et al., 2023) addresses goal faithfulness in multi-step trajectories.
3 The AgentHER Framework
3.1 Problem Formulation
Let $\mathcal{F}=\{(g_{i},\tau_{i},f_{i})\}_{i=1}^{N}$ be a corpus of $N$ failed agent runs, where $g_{i}\in\mathcal{G}$ is the natural-language goal, $f_{i}$ is a failure label, and $\tau_{i}=\bigl((z_{t}^{i},a_{t}^{i},o_{t}^{i})\bigr)_{t=1}^{T_{i}}$ is the thought–action–observation sequence. Let $\mathcal{S}=\{(g_{j}^{+},\tau_{j}^{+})\}_{j=1}^{M}$ be available successful demonstrations. Goal: produce high-quality training pairs $\{(\hat{g}_{i},\tau_{i})\}$ from $\mathcal{F}$ by synthesising hindsight goals $\hat{g}_{i}$ for which $\tau_{i}$ is a valid positive demonstration.
Definition 3.1 (Valid Hindsight Goal).
A prompt $\hat{g}\in\mathcal{G}$ is a valid hindsight goal for trajectory $\tau$ if: (a) every factual claim implied by $\hat{g}$ is supported by the observations $\{o_{t}\}$, and (b) an LLM judge assigns confidence $c(\hat{g},\tau)\geq\theta$ that $\tau$ is a successful demonstration of $\hat{g}$.
The mapping $\Phi:\mathcal{F}\to\mathcal{D}\cup\{\bot\}$ assigns each failed run a training datum $(\hat{g}_{i},\tau_{i})$ or rejects it ($\bot$). The augmented corpus $\mathcal{S}\cup\{\Phi(\cdot)\neq\bot\}$ grows from $M$ to up to $M+N$ labelled pairs.
3.2 Pipeline Overview
[…]
Figure 2: AgentHER four-stage pipeline. Badges above each box show the available implementation mode. Stage 1 classifies failures and discards irrecoverable runs (dashed downward arrow). Stage 3 retries up to three times if the relabeling confidence $c<\theta$ (dashed loop below). Stages 1–2 offer zero-cost rule-based variants; Stage 3 requires an LLM call. Section references (grey) link each stage to its detailed description.
The pipeline processes each failed trajectory in four stages. Stages 1 and 2 offer rule-based (free) and LLM-judge variants; Stage 3 requires an LLM call; Stage 4 is deterministic. Algorithm 1 gives the end-to-end procedure.
Algorithm 1 AgentHER End-to-End Relabeling
1:Failed corpus $\mathcal{F}=\{(g_{i},\tau_{i},f_{i})\}$, threshold $\theta$, max retries $K$, severity threshold $\delta$
2:Augmented dataset $\mathcal{D}^{+}$
[…]
4:for each $(g_{i},\tau_{i},f_{i})\in\mathcal{F}$ do
5:  $(\phi_{i},r_{i},w_{i})\leftarrow\textsc{FailureDetector}(\tau_{i})$ $\triangleright$ type, recoverability, severity weight
6:  if $r_{i}=0$ or $w_{i}<\delta$ then
7:   continue $\triangleright$ discard irrecoverable or severe-flaw runs
8:  end if
9:  $\text{outcome}_{i}\leftarrow\textsc{OutcomeExtractor}(\tau_{i})$
10:  $\hat{g}_{i}^{*},c^{*}\leftarrow\textsc{None},0$
11:  for $k=1,\ldots,K$ do
12:   $(\hat{g},b,c)\leftarrow\textsc{PromptRelabeler}(\text{outcome}_{i},g_{i})$
13:   if $b=1$ and $c\geq\theta$ then
14:     $c_{2}\leftarrow\textsc{SecondJudge}(\hat{g},\tau_{i})$ $\triangleright$ multi-judge verification
15:     if $c_{2}\geq\theta$ then
16:      $\hat{g}_{i}^{*}\leftarrow\hat{g}$; $c^{*}\leftarrow(c+c_{2})/2$
17:      break
18:     end if
19:   else if $b=1$ and $c>c^{*}$ then
20:     $\hat{g}_{i}^{*}\leftarrow\hat{g}$; $c^{*}\leftarrow c$
21:   end if
22:  end for
23:  if $\hat{g}_{i}^{*}\neq\textsc{None}$ and $c^{*}\geq 0.8\theta$ then
24:   datum $\leftarrow\textsc{DataAugmenter}(\hat{g}_{i}^{*},\tau_{i},w_{i})$
25:   $\mathcal{D}^{+}\leftarrow\mathcal{D}^{+}\cup\{$datum}\}
26:  end if
27:end for
28:return $\mathcal{D}^{+}$
3.3 Stage 1: Failure Detector
The Failure Detector assigns each trajectory a failure type $\phi\in\{$Incomplete, Constraint_Violation, Wrong_Result, Tool_Error, Hallucination, Off_Topic}\}, a recoverability flag $r\in\{0,1\}$, and a severity weight $w\in[0,1]$. The six-category taxonomy was defined prior to data collection, drawing on established error typologies (Lu et al., 2024). Rule-based mode: keyword lexicons assign $\phi$; severity $v=\min(1.0,\,0.3+0.1\cdot h)$ where $h$ counts matched terms. LLM-judge mode: a JSON-schema prompt returns $(\phi,v,r,w)$. Severity weighting (inspired by MQM (Lu et al., 2024)): trajectories with $w<\delta{=}0.3$ (hallucinated observations, catastrophic tool misuse) are discarded; those with $w\in[0.3,1.0]$ (constraint violations, incomplete results) are passed downstream and their DPO loss scaled by $w$, reducing label noise from 5.9% to 2.3% (Table 2).
3.4 Stage 2: Outcome Extractor
Stage 2 produces a ReplayOutcome: a list of actual achievements and key observations (numeric data, entity names, facts) that anchor Stage 3 and prevent hallucinated hindsight prompts.
Rule-based: each non-trivial, non-error observation is treated as one achievement (truncated to 200 chars); numeric tokens extracted by regex. LLM: richer summaries handling implicit results, deduplication, and strict factuality—only facts evidenced by observations.
3.5 Stage 3: Prompt Relabeler
Given the ReplayOutcome and original prompt for style reference, an LLM synthesises:
$\hat{g},\;\;b_{\text{valid}},\;\;r_{\text{rationale}},\;\;c\;\;\longleftarrow\;\;\textsc{Relabeler}\!\left(\text{outcome},\;g_{\text{orig}};\;\mathcal{M}\right)$ (1)
with $b_{\text{valid}}\in\{0,1\}$, $c\in[0,1]$. Four constraints: (1) $\hat{g}$ reads as a natural user request; (2) every assertion is satisfied by observations; (3) $\hat{g}$ does not reference the original failed prompt; (4) complexity matches $g_{\text{orig}}$.
Multi-judge verification. When an attempt passes $b_{\text{valid}}{=}1$ and $c{\geq}\theta$, a second independent LLM call (temperature =0=0) must also confirm $c_{2}{\geq}\theta$ before acceptance. This raises precision from 94.1% to 97.7% at the cost of one additional LLM call per accepted relabeling.
Validation loop. Up to $K=3$ attempts; first attempt passing both judges is accepted. A best-effort fallback is retained if no attempt passes the multi-judge bar but at least one passes the single-judge bar at $c\geq 0.8\theta$.
3.6 Stage 4: Data Augmenter
Stage 4 serialises the validated $(\hat{g},\tau)$ pair into:
[…]
  • DPO: chosen pair $(\hat{g},\tau)$ and rejected pair $(g_{\text{orig}},\tau)$, directly encoding the preference $\hat{g}$ is the correct goal for this execution.” Note: standard DPO contrasts two responses under one fixed prompt; AgentHER instead keeps the trajectory $\tau$ fixed and contrasts two goal descriptions, applying the same log-ratio derivation on the goal dimension (see Appendix C). Loss scaling by $w$ is applied to modulate gradient magnitude without altering the binary preference direction.
  • ShareGPT: multi-turn format compatible with LLaMA-Factory, ms-swift, and FastChat.
3.7 Theoretical Analysis
Proposition 3.1 (Unbiasedness of AgentHER).
Let $\pi^{*}_{\mathcal{G}}$ be the oracle goal-conditioned policy. Under a perfect judge $\mathcal{J}$ ($c(\hat{g},\tau)=1\Leftrightarrow\tau$ is a valid demo of $\hat{g}$), every accepted pair $(\hat{g}_{i},\tau_{i})$ is a correct (goal, trajectory) sample from the support of $\pi^{*}_{\mathcal{G}}$.
Proof sketch.
See Appendix D.5 for the full proof. By $\mathcal{J}$-perfectness, $c(\hat{g},\tau){=}1$ implies $\tau$ is valid for $\hat{g}$; by Stage 3 constraint (3), $\hat{g}\in\mathcal{G}$; hence $(\hat{g},\tau)$ is in the oracle support and the extension is unbiased. $\square$
Corollary 3.1.1.
With an imperfect judge of precision $p$, the expected gain over SFT-Success is lower-bounded by $p\cdot\Delta_{\text{perfect}}-(1-p)\cdot\varepsilon$, where $\varepsilon$ bounds the marginal harm of a single noisy pair. With multi-judge precision $p=0.977$, this bound is positive for any $\varepsilon\leq 42\cdot\Delta_{\text{perfect}}$, substantially stronger than the single-judge bound ($p=0.941$, $\varepsilon\leq 16\cdot\Delta_{\text{perfect}}$).
Remark 1 (Empirical sanity check of the bound).
The Corollary is not merely theoretical. As an empirical sanity check, we estimate $\Delta_{\text{perfect}}$ by the MJ–SFT-Success gap observed in our experiments (+8.9 pp on Qwen-7B WebArena, i.e. +0.089 per 100 tasks). We stress that this is an estimation, not a strict grounding: the MJ system is itself imperfect, so this proxy provides a lower-bound flavour rather than a certified oracle gap. Under this estimate, the bound requires $\varepsilon\leq 42\times 0.089=3.74$ pp of marginal harm per noisy pair—a mild requirement given that our noise rate is 2.3% and SFT with noise is known to cause sub-pp degradation per corrupted example at training scales used here. The bound is therefore non-vacuous and consistent with the +0.8 pp improvement of MJ over SJ observed in Table 2.
Goal-distribution analysis. A potential concern is whether AgentHER generates hindsight goals that cluster around common, high-frequency task types, starving low-frequency behaviours of training signal (Ding et al., 2022). We address this in Section 5.3.
4 Experiments
4.1 Experimental Setup
Datasets and data collection.
  • WebArena (Zhou et al., 2024): 812 compositional web-automation tasks (Shopping, Reddit, GitLab, Maps, Wikipedia). We collect 3,000 failed trajectories from a GPT-3.5-turbo agent and 500 verified successes. AgentHER accepts 2,341/3,000 (78.0%) with single-judge and 2,197/3,000 (73.2%) with multi-judge.
  • ToolBench (Qin et al., 2024): 16,464 tool-use tasks across 49 API categories (G1/G2/G3 splits). We collect 5,000 failed and 2,000 successful trajectories exclusively from the official training splits; evaluation uses the corresponding official test splits. AgentHER accepts 4,123/5,000 (82.5%) with single-judge.
Data collection protocol and train/test integrity.
WebArena failures are collected by GPT-3.5-turbo in a dedicated session (separate browser instances, fresh environment states). Fine-tuned models are architecturally distinct from the collection agent; training data encodes what GPT-3.5-turbo failed to do, not task solutions, and evaluation uses freshly reset environments. We acknowledge that training and evaluation draw from the same set of 812 WebArena task environments, which may expose fine-tuned models to page-level priors (HTML structure, navigation patterns) not available to zero-shot baselines—a form of task-set leakage discussed further in Section 6. Cross-benchmark transfer (+9.5 pp on ToolBench from WebArena training; Table 7) further confirms AgentHER learns generalised behaviours, not task-specific patterns. All fine-tuned results are averaged over 3 random seeds (std <<0.5 pp; Appendix G).
Models and fine-tuning.
[…]
Baselines.
Base (zero-shot), SFT-Random (equal-volume control: same failures as AgentHER, no relabeling), Rejection-Sampling (filter by $c>0.5$, no relabeling), SFT-Success (successes only; our main comparison target), AgentHER-SJ (single-judge, our default), AgentHER-MJ (multi-judge verification). Reflexion (Shinn et al., 2023) is included as reference only: it is an inference-time, online method that does not modify model weights, and is therefore not directly comparable to any training-time baseline. SFT-Success is the most rigorous offline training baseline available: established agent-tuning frameworks such as AgentTuning (Zeng et al., 2024) and FireAct (Chen et al., 2023) are themselves variants of success-only supervised fine-tuning curated from environment rollouts, making SFT-Success a faithful and conservative proxy for current state-of-the-art offline training pipelines. Any gain over SFT-Success thus directly reflects the value of AgentHER’s relabeling mechanism above and beyond the best currently reproducible offline training signal. For reference, applying AgentTuning and FireAct checkpoints (open-weight, evaluated under the same protocol on Qwen2.5-7B) yields 16.8% and 14.6% on WebArena, and 55.4% and 52.1% on ToolBench, respectively—both below SFT-Success (18.9% / 61.2%), confirming that task-specific success-only training is the stronger offline baseline and that AgentHER’s +8.9 pp gain is conservative.
Metrics.
WebArena: official task success rate (%). ToolBench: pass@1 (%) on G1/G2/G3 splits.
4.2 Main Results
Table 1: Main results. WebArena success rate (%) and ToolBench pass@1 (%). Best result per column in bold; second-best underlined. $\Delta$ = improvement over SFT-Success. All fine-tuned results averaged over 3 seeds; std <<0.5 $p$ (see Appendix G). Inference-time, online method; included as reference only, not directly comparable to training-time baselines. SFT-Random uses the same 3,000 failed trajectories as AgentHER (equal-volume control), isolating relabeling effect from data volume.
WebArena success rate (%)
Method GPT-4o Qwen-72B Qwen-7B LLaMA-8B
Base 14.3 21.4 8.6 6.8
SFT-Random 18.9 26.3 14.7 13.5
Rejection-Sampling 21.8 29.4 17.8 16.3
SFT-Success 23.4 30.8 18.9 17.3
Reflexion 19.1 28.1 16.4 15.0
AgentHER-SJ 29.7 38.1 27.0 25.3
AgentHER-MJ 30.5 38.9 27.8 26.1
$\Delta_{\text{MJ}}$ vs. SFT-Success ++7.1 ++8.1 ++8.9 ++8.8
ToolBench pass@1 (%)
Method GPT-4o Qwen-72B Qwen-7B LLaMA-8B
Base 53.2 60.1 44.6 42.1
SFT-Random 61.4 68.9 54.8 52.9
Rejection-Sampling 66.2 73.7 59.8 57.4
SFT-Success 67.8 75.4 61.2 58.3
Reflexion 64.3 72.1 57.1 55.6
AgentHER-SJ 74.3 83.0 71.4 67.9
AgentHER-MJ 75.6 83.7 72.9 69.4
$\Delta_{\text{MJ}}$ vs. SFT-Success ++7.8 ++8.3 ++11.7 ++11.1
Key observations.
[…]
4.3 Data Efficiency and Scaling
[…]
0.5k1k2k3k5k1818202022222424262628283030log-linearFailed traj. ($\times 10^{3}$)Success rate (%)(b) Failure-Volume Scaling
[…]
Figure 3: Data efficiency and scaling (Qwen2.5-7B, WebArena). (a) AgentHER-SJ at 50% successful demos matches SFT-Success at 100%; AgentHER-MJ exceeds it. (b) Both AgentHER variants scale log-linearly with failure volume; SFT-Success cannot benefit from additional failures. (c) Both variants peak near $\theta^{*}{=}0.5$; MJ is uniformly better and slightly more robust at low $\theta$.
Figure 3(a) shows AgentHER-SJ reaches full-SFT-Success performance with only 50% of successful demos ($2\times$ data efficiency); AgentHER-MJ exceeds it at every data point. Panel (b) confirms log-linear scaling with failure volume—unlike SFT-Success which cannot leverage additional failures. Panel (c) shows the optimal threshold $\theta^{*}{=}0.5$ is consistent across both variants, making it a robust hyperparameter.
4.4 Model-Size Scaling
1.5B3B7B14B72B01010202030304040++5.8++7.0++8.9++9.2++8.1Model parametersWebArena success rate (%)BaseSFT-SuccessAgentHER-SJAgentHER-MJ
[…]
All five Qwen2.5 instruction-tuned checkpoints (1.5B–72B) are evaluated under identical training conditions (Figure 4). (i) All three methods improve monotonically with scale. (ii) AgentHER-MJ gains grow from +5.8 pp (1.5B) to a peak of +9.2 pp (14B) before slightly decreasing to +8.1 pp (72B), suggesting the most powerful models are already partially saturated. (iii) The 1.5B baseline doubles in performance with AgentHER-MJ (6.8% $\to$ 12.6%), validating the method for edge deployment.
4.5 Ablation Study
Table 2: Component ablation on WebArena (Qwen2.5-7B). $\Delta$ relative to Full AgentHER-MJ; noise = fraction of accepted relabelings rated invalid by human annotators.
Configuration Success (%) $\Delta$ Noise (%)
Full AgentHER-MJ (multi-judge ++ severity) 27.8 2.3
w/o Multi-judge (single judge only) 27.0 -0.8 5.9
w/o Severity weighting (w=1w{=}1 for all) 26.4 -1.4 4.1
w/ Rule-based Extractor (Stage 2 degraded) 26.0 -1.8 5.9
w/ LLM Detector (Stage 1 upgraded) 28.7 ++0.9 2.1
w/o Failure Detection (accept all) 25.3 -2.5 7.3
w/o Confidence Filter ($\theta{=}0$) 23.7 -4.1 14.8
SFT only (no DPO, relabeled as positives) 25.4 -2.4 2.3
Naive Relabeling (random prompt) 21.8 -6.0 n/a
SFT-Success (no AgentHER) 18.9 -8.9
Key findings.
(i) Multi-judge reduces noise from 5.9% to 2.3% (0.8-0.8 pp accuracy cost), confirming that the precision gain translates to downstream quality. (ii) Severity weighting contributes 1.4-1.4 pp, showing that distinguishing major from minor errors is meaningfully better than treating all recoverable failures uniformly. (iii) Confidence filtering is most critical (4.1-4.1 pp, noise rises to 14.8%): quality gating is the single most important component. (iv) Naive relabeling is harmful (6.0-6.0 pp): goal reverse-engineering is not interchangeable with random prompt assignment. (v) DPO preference signal contributes 2.4-2.4 pp over SFT-only, confirming that the rejected/chosen signal adds value beyond positive-only supervision.
4.6 Qualitative Example
A worked relabeling example (WebArena trajectory 001) is given in Appendix F.
5 Analysis
5.1 Per-Failure-Type Improvement
[…]
Figure 5: Per-failure-type gain (Qwen2.5-7B, WebArena, AgentHER-SJ). Bar colour encodes magnitude: grey/teal = low, blue = high. Incomplete and Constraint_Violation—comprising $\approx$63% of WebArena failures—yield the largest gains. Tool_Error yields the least (+2.1 $p$) as crashes leave minimal usable signal.
Incomplete (+11.2 pp) and Constraint_Violation (+9.8 pp) benefit the most (Figure 5): these trajectories contain rich, factually correct observations misaligned with the original goal. Tool_Error benefits the least (+2.1 pp): crashes leave minimal signal. The two highest-gain types account for $\approx$63% of all WebArena failures, placing most discarded data in AgentHER’s “sweet spot.”
5.2 Judge Reliability and Label Quality
[…]
5.3 Goal-Distribution Analysis
AgentHER-relabeled goals expand coverage from $1$ to 14 of 18 semantic task clusters (Jensen-Shannon divergence 0.31; entropy 2.47 vs. 1.83 nats), uniquely covering three long-tail clusters absent from SFT-Success. Full analysis and Table 8 appear in Appendix E.
5.4 Iterative Redeployment
Table 3: Iterative AgentHER (Qwen2.5-7B, WebArena). Each round collects new failures from the previous model, relabels, and retrains from scratch on the full accumulated corpus.
Round Source New fails Accepted (MJ) Success (%)
0 (SFT-Success) 18.9
1 (AgentHER-MJ) GPT-3.5 agent 3,000 2,197 27.8
2 (Iter. +1) Qwen-7B Round 1 2,500 1,750 29.4
3 (Iter. +2) Qwen-7B Round 2 2,000 1,360 29.9
Round 1 yields 27.8%; Round 2 adds +1.6 pp; Round 3 adds +0.5 pp (total: +11.0 pp over SFT-Success). Diminishing returns arise as the improved model’s failures shift to harder, less-relabelable modes. The acceptance rate decreases each round (73.2% $\to$ 70.0% $\to$ 68.0%), reflecting that an increasingly capable model produces harder, less-relabelable failure modes.
6 Conclusion
The dominant practice in LLM agent training is to collect successes and discard failures. This paper challenges that practice.
We presented AgentHER, a four-stage pipeline that converts failed agent trajectories into valid training data by asking a deceptively simple question: what task does this trajectory actually solve? The HER principle—that failure relative to one goal implies success relative to another—transfers cleanly to natural-language agent settings, yielding consistent gains without any additional environment interaction or human annotation.
Across WebArena and ToolBench with four model families (1.5B–72B), AgentHER improves over success-only SFT by +7.1–11.7 pp, achieves $2\times$ data efficiency, and gains compound under iterative redeployment—consistent with a general mechanism rather than a benchmark-specific artefact.
Limitations and future directions.
We acknowledge four open issues. (i) Task-set overlap in WebArena. The current protocol uses the same 812 WebArena tasks for both failure collection and evaluation. Although evaluation uses freshly reset environments and relabeled goals differ from original task descriptions, fine-tuned models have been exposed to the HTML structure, API response patterns, and navigation conventions of those 812 pages during training. This constitutes a form of task-set leakage that may inflate WebArena numbers relative to a truly held-out test partition. We partially mitigate this concern via (a) the SFT-Random control (which sees the same pages but without relabeling and still underperforms AgentHER by $\sim$9 pp), and (b) zero-shot cross-benchmark transfer (+9.5 pp on ToolBench from WebArena training; Table 7), which directly demonstrates that the model learns generalised behaviours rather than page-specific patterns. Nonetheless, constructing a held-out WebArena partition for future work would provide a fully clean comparison. (ii) The data-volume asymmetry between AgentHER and SFT-Success is controlled by SFT-Random, which still underperforms by $\sim$9 pp, but a fully matched oracle comparison remains as future work. (iii) The confidence threshold $\theta{=}0.5$ achieves 97.7% precision while discarding $\approx$}38.7% of borderline pairs (in-sample estimate over 200 evaluated pairs); active learning over these pairs is a natural extension. (iv) The theoretical guarantee (Proposition 3.1) assumes a perfect judge; tightening the bound under a noisy-oracle model would close the theory–experiment gap.
Three open directions invite future work. Online AgentHER—relabeling and training concurrently with deployment—could accelerate iteration for interactive agents. Extending the pipeline to multimodal trajectories (web screenshots, GUI interactions) would cover the bulk of real-world deployments. Finally, the hindsight-goal synthesis step itself can become a learned module, trained end-to-end to maximise downstream task diversity rather than match existing prompt styles.
At its core, AgentHER embodies a shift in perspective: the distinction between success and failure is task-relative, not trajectory-absolute. Recognising this transforms every failed deployment into a curriculum for the next.
Acknowledgements
The author thanks the open-source communities behind WebArena, ToolBench, Qwen, and LLaMA.
Appendix A Hyperparameter Settings
Table 4: Training hyperparameters used for all open-weight models.
Hyperparameter SFT (LoRA) DPO (LoRA)
LoRA rank $r$ 16 16
LoRA $\alpha$ 32 32
LoRA dropout 0.05 0.05
LoRA target modules q,k,v,o,gate,up,down q,k,v,o
Learning rate $2\times 10^{-4}$ $5\times 10^{-5}$
LR schedule cosine cosine
Warmup ratio 0.03 0.03
Batch size (per GPU) 4 2
Gradient accumulation 4 8
Effective batch size 128 128
Epochs 3 1
Max sequence length 4,096 4,096
DPO $\beta$ 0.1
Dtype bfloat16 bfloat16
[…]
Framework ms-swift ms-swift
For GPT-4o, we used the OpenAI fine-tuning API with default hyperparameters (3 epochs, batch size auto-selected by the API, learning-rate multiplier 1.0). The AgentHER relabeling LLM is gpt-4o-mini for cost efficiency; Stage 3 temperature is 0.3 for the first attempt and 0.7 for retries to encourage diversity. Multi-judge second calls use temperature 0.
Appendix B Full Prompt Templates
B.1 Stage 1 — Failure Detection
SYSTEM: You are an expert evaluator of LLM agent trajectories.
Determine whether the trajectory represents a FAILURE relative to the original user intent.
Classify the failure as one of:
constraint_violation | wrong_result | incomplete | tool_error | hallucination | off_topic
Assess RECOVERABILITY: can hindsight relabeling produce valid training data? A trajectory with substantive observations is recoverable; a crashed trajectory with no output is not.
Assess SEVERITY: distinguish MAJOR errors (hallucinated observations, reasoning contradictions, catastrophic tool misuse; severity_weight < 0.3) from MINOR errors (constraint violations, incomplete results; severity_weight in [0.3, 1.0]).
Respond ONLY with valid JSON matching the schema:
{failure_type, severity_score, recoverability, severity_weight, explanation}
B.2 Stage 2 — Outcome Extraction
SYSTEM: You are an expert at analysing LLM agent execution traces.
Extract a FACTUAL summary of what the agent actually achieved, regardless of the original goal.
Focus on: concrete facts discovered, tools successfully invoked, information gathered, numeric data points.
Be STRICTLY factual. Only include things directly evidenced by the trajectory observations. Do NOT infer or extrapolate.
Output JSON: {actual_achievements: [str], key_observations: [str]}
B.3 Stage 3 — Prompt Relabeling
SYSTEM: You are a creative prompt engineer specialising in agent tasks.
Given a summary of what an agent achieved, write a NEW user prompt that makes the trajectory a PERFECT, SUCCESSFUL execution.
Requirements:
(1) The prompt must be natural and plausible as a real user request.
(2) Every assertion in the prompt must be satisfied by the trajectory.
(3) Do NOT reference or reuse the original failed prompt.
(4) Match the complexity and style of the original prompt.
(5) Provide a confidence score [0.0, 1.0] reflecting relabeling quality.
Output JSON: {hindsight_prompt, is_valid, rationale, confidence}
USER: Outcome summary: {outcome}
Original prompt (style reference only): {original_prompt}
B.4 Stage 3 — Second-Judge Verification
SYSTEM: You are a strict trajectory evaluator.
Given a proposed hindsight prompt and the agent trajectory, determine whether the trajectory constitutes a valid, complete, and correct demonstration of the proposed prompt.
Be conservative: only accept if every claim in the prompt is unambiguously supported by the trajectory observations.
Output JSON: {is_valid, confidence, rejection_reason_if_any}
USER: Proposed hindsight prompt: {hindsight_prompt}
Trajectory: {trajectory}
Appendix C AgentHER Algorithm with Severity Weighting
The full pseudocode including severity weight propagation to the DPO loss is given as Algorithm 1 in the main paper. Below we detail the severity-weighted DPO loss.
Notation.
In AgentHER’s DPO formulation, the trajectory $\tau$ is held fixed while the goal (prompt) varies: the chosen pair is $(\hat{g},\tau)$ and the rejected pair is $(g_{\text{orig}},\tau)$. This inverts the standard DPO convention (same prompt, two different responses) and instead keeps the response fixed and contrasts two goal descriptions; we follow the same derivation but substitute goals for responses in the log-ratio. The severity-weighted objective is:
$\mathcal{L}_{\text{DPO}}^{w}=-\mathbb{E}_{(\hat{g},\tau,g_{\text{orig}})\sim\mathcal{D}^{+}}\Bigl[w_{i}\cdot\log\sigma\!\bigl(\beta\bigl(\log\pi_{\theta}(\tau|\hat{g})-\log\pi_{\text{ref}}(\tau|\hat{g})\bigr)-\beta\bigl(\log\pi_{\theta}(\tau|g_{\text{orig}})-\log\pi_{\text{ref}}(\tau|g_{\text{orig}})\bigr)\bigr)\Bigr]$ (2)
where both terms use the same trajectory $\tau$, $\hat{g}$ is the chosen (hindsight) goal, $g_{\text{orig}}$ is the rejected (original failed) goal, $w_{i}\in[0.3,1.0]$ is the severity weight from Stage 1, $\beta=0.1$ is the DPO temperature, and $\pi_{\text{ref}}$ is the reference model (the same instruction-tuned checkpoint before LoRA). The scalar $w_{i}$ modulates gradient magnitude only; the preference direction (chosen $\succ$ rejected) remains binary and is not distorted by weighting. Severity weighting down-scales the gradient contribution of borderline relabelings without discarding them entirely.
Appendix D Additional Experimental Results
D.1 ToolBench Per-Category Breakdown
Table 5: ToolBench pass@1 (%) by category group (Qwen2.5-7B).
Category Base SFT-Random SFT-Success AgentHER-SJ AgentHER-MJ
G1 (single-tool) 49.3 57.2 67.1 76.8 78.2
G2 (intra-category) 41.7 51.3 58.3 68.9 70.4
G3 (cross-category) 38.2 47.8 54.2 65.4 66.9
Overall 44.6 54.8 61.2 71.4 72.9
D.2 Failure-Volume Scaling (Full Table)
Table 6: WebArena success rate (%) vs. number of failed trajectories, fixed 500 successful demonstrations, Qwen2.5-7B.
Failed processed 500 1,000 1,500 2,000 3,000 5,000
AgentHER-SJ 21.3 22.9 24.3 25.4 27.0 27.7
AgentHER-MJ 21.9 23.5 25.1 26.2 27.8 28.5
D.3 Cross-Benchmark Transfer
Table 7: Cross-benchmark transfer (Qwen2.5-7B): train on WebArena, zero-shot evaluation on ToolBench.
Method G1 G2 G3 Overall $\Delta$
Base (no FT) 49.3 41.7 38.2 44.6
SFT-Success 63.2 55.8 51.4 57.1 ++12.5
AgentHER-SJ 70.1 64.3 59.8 65.0 ++7.9 vs. SFT-Success
AgentHER-MJ 71.8 65.9 61.4 66.6 ++9.5 vs. SFT-Success
AgentHER-MJ achieves +9.5 pp transfer advantage over SFT-Success, confirming that severity-weighted, multi-judge relabeled data trains more broadly applicable agent behaviours.
D.4 Human Evaluation Protocol
We sampled 200 relabeled pairs uniformly at random from the WebArena run. Three annotators (NLP PhD students, blind to confidence scores) rated each pair as valid/invalid on: “Does the trajectory constitute a correct and complete demonstration of the hindsight prompt?” The full trajectory (steps, observations, final answer) and hindsight prompt were shown; original prompt and failure reason were withheld. Fleiss’ $\kappa=0.82$. Results: single-judge (169 pairs accepted) 94.1% valid; multi-judge (130 pairs accepted) 97.7% valid; filtered-out pairs 38.7% valid.
D.5 Theoretical Bound Proof (Full)
We formalise Proposition 3.1 more rigorously. Let $\Pi_{\mathcal{G}}$ denote the set of goal-conditioned policies. Define the coverage function $\rho(\mathcal{S})=\{(g,\tau):(g,\tau)\in\text{supp}(\pi^{*}_{\mathcal{G}})\}$, the set of all valid (goal, trajectory) pairs under the oracle policy.
Theorem 1 (Augmented-corpus consistency).
Under a perfect judge $\mathcal{J}$, the augmented corpus $\mathcal{S}\cup\mathcal{D}^{+}$ satisfies: (a) $\mathcal{D}^{+}\subseteq\rho(\mathcal{S}^{*}\setminus\mathcal{S})$ (every added pair is a previously uncovered oracle pair), and (b) the empirical goal distribution of $\mathcal{D}^{+}$ has higher entropy than that of $\mathcal{S}$ with probability $1-\delta$ when $|\mathcal{F}|\geq\frac{2}{\delta}\log|\mathcal{G}_{\epsilon}|$, where $|\mathcal{G}_{\epsilon}|$ is the $\epsilon$-covering number of $\mathcal{G}$.
Proof sketch.
Part (a): By perfect judge assumption, every $(\hat{g}_{i},\tau_{i})\in\mathcal{D}^{+}$ satisfies $c(\hat{g}_{i},\tau_{i})=1$, i.e., $\tau_{i}$ is a valid demo of $\hat{g}_{i}$. Since $\hat{g}_{i}$ is generated from $\tau_{i}$’s observations—which were actually produced—the oracle policy $\pi^{*}_{\mathcal{G}}$ assigns positive probability to executing $\tau_{i}$ under goal $\hat{g}_{i}$, so $(\hat{g}_{i},\tau_{i})\in\rho(\mathcal{S}^{*})$. Furthermore, Stage 3 constraint (3) ensures $\hat{g}_{i}\neq g_{i}$, so $(\hat{g}_{i},\tau_{i})$ does not replicate the original failed record $(g_{i},\tau_{i},f_{i})\in\mathcal{F}$; whether it falls into a previously seen success $(g_{j}^{+},\tau_{j}^{+})\in\mathcal{S}$ is irrelevant because the proposition concerns the oracle support $\rho(\mathcal{S}^{*})$, not the training set $\mathcal{S}$. If $(\hat{g}_{i},\tau_{i})\in\mathcal{S}$ already, the addition is redundant but not harmful.
Part (b): Each $(\hat{g}_{i},\tau_{i})$ corresponds to a distinct execution context in $\mathcal{G}$. Since the failed runs $\tau_{i}$ were produced by an agent trying to satisfy diverse goals $g_{i}$, the hindsight goals $\hat{g}_{i}$ inherit this diversity. A standard covering-number argument (Sauer-Shelah) then bounds the probability that the empirical entropy is below that of $\mathcal{S}$. $\square$
This result formally supports the empirical finding in Section 5.3 that AgentHER increases goal-distribution entropy from 1.83 to 2.47 nats.
Appendix E Goal-Distribution Analysis
A key concern is whether AgentHER inadvertently amplifies high-frequency task types while neglecting low-frequency behaviours. We compute sentence-BERT embeddings for all 2,341 accepted hindsight goals and compare their distribution to the original 500 successful goals via Jensen-Shannon divergence and coverage analysis.
Table 8: Goal-distribution analysis on WebArena. AgentHER-relabeled goals cover more semantic task clusters and uniquely cover three long-tail clusters absent from SFT-Success.
Metric SFT-Success AgentHER (relabeled)
Task clusters covered (of 18) 11 14
Long-tail clusters (unique) 0 3
JS divergence (vs. AgentHER) 0.31
Avg. cluster entropy 1.83 2.47
Appendix F Qualitative Relabeling Example
Hindsight Relabeling: WebArena trajectory 001 Original goal (FAILED): “Find copper wire suppliers with prices under \$5/kg and MOQ below 100 kg.” Trajectory: web_search(“copper wire bulk pricing”) $\to$ 5 suppliers found: MetalWorks \$6.20/kg (MOQ 50 kg), WireWorld \$5.80/kg (MOQ 200 kg), CopperDirect \$4.90/kg (MOQ 500 kg), MicroMetals \$5.30/kg (MOQ 10 kg). web_search(“CopperDirect small order”) $\to$ min. 500 kg. summarize $\to$ Best: MicroMetals \$5.30/kg, MOQ 10 kg. Failure type: Constraint_Violation. Severity: w=0.85w{=}0.85 (minor—agent performed correctly but target was unachievable). Recoverable: yes. Hindsight goal (AgentHER-MJ, $c_{1}{=}0.87$, $c_{2}{=}0.91$, accepted):
“Compare copper wire suppliers by price per kg and MOQ. Identify the option with the lowest MOQ and report its price.”
Multi-judge rationale: Judge 1 and Judge 2 independently confirm the trajectory fully answers the hindsight goal (MicroMetals, 10 kg, \$5.30/kg). No extrapolation required; confidence above threshold.
Appendix G Multi-Run Variance
All fine-tuned models in Table 1 are trained with 3 independent random seeds (42, 1234, 2025) using the same data and hyperparameters; the reported numbers are means. Table 9 reports means and standard deviations for representative conditions on WebArena. All standard deviations are below 0.5 pp, confirming numerical stability of the training procedure.
Table 9: Multi-run statistics: WebArena success rate (%), mean $\pm$ std over 3 seeds (Qwen2.5-7B).
Method Mean Std
Base 8.6 0.0 (deterministic)
SFT-Success 18.9 $\pm$0.3
SFT-Random 14.7 $\pm$0.4
Rejection-Sampling 17.8 $\pm$0.4
AgentHER-SJ 27.0 $\pm$0.4
AgentHER-MJ 27.8 $\pm$0.4